OpenStudy (anonymous):

Find the equations of the tangent lines at the point the curve crosses itself: x=t^3-6t y= t^2 I came up with the following

OpenStudy (anonymous):

\[y= \pm \frac{ 12 }{ \sqrt{6} }x\]

OpenStudy (anonymous):

what did you do?

OpenStudy (anonymous):

\[t _{1}^{3}-6t _{1}=t _{2}^{3}-6t _{2}.....AND....t _{1}^{2}=t _{2}^{2}\]

OpenStudy (anonymous):

correct but you also need to assume that \[t_1 \neq t_2\]

OpenStudy (anonymous):

\[t _{1}neqt _{2}\]

OpenStudy (anonymous):

I did left that out...So t1=-t2

OpenStudy (anonymous):

and t2=-t1

OpenStudy (anonymous):

you only need t1 = -t2. There is not need for the other equation since you just move the negative to the other side.

OpenStudy (anonymous):

so the two values of the parameter is -sqrt(6) and sqrt(6) yes?

OpenStudy (anonymous):

I substituted back into the first equation and ended up with \[2t _{1}(t _{1}^{2}-6)=0 ....thus t=0 ..and..t=\sqrt{6}\]

OpenStudy (anonymous):

The lines crosses itself at (0,6)

OpenStudy (anonymous):

if t1= 0, then t2 = 0 but we already assume t1 =/= t2. also, t^2 - 6 = 0 t^2 = 6 t =+/- sqrt(6)

OpenStudy (anonymous):

\[\frac{ 2t }{ 3t ^{2}-6 }\]

OpenStudy (anonymous):

or \[\frac{ 2\sqrt{6} }{ 12 }\]

OpenStudy (anonymous):

that's one slope. What about the other one?

OpenStudy (anonymous):

Negative

OpenStudy (anonymous):

right. Btw, simply the value of slope further.

OpenStudy (anonymous):

\[y=\frac{ \sqrt{6} }{ 6 }x-\sqrt{6}\]

OpenStudy (anonymous):

and\[y= -\frac{ \sqrt{6} }{ 6 }x+\sqrt{6}\]

OpenStudy (anonymous):

it's \[y=\frac{ \sqrt{6} }{ 6 }x+\sqrt{6}\]

OpenStudy (anonymous):

the point is (0,6) so answer is y = +/- sqrt(6)x/6 + sqrt(6)

OpenStudy (anonymous):

\[\frac{ 2\sqrt{6} }{ 12 }(x-6)=\frac{ 2\sqrt{6} }{ 12 }-\frac{ 12\sqrt{6} }{ 12 }\]

OpenStudy (anonymous):

again the point is (0,6). NOT (6,0)

OpenStudy (anonymous):

My bad...

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

the sqrt (6) remains positive for both equations?

OpenStudy (anonymous):

i made a mistake it's y = +/- sqrt(6)/6x + 6

OpenStudy (anonymous):

yes it's positive for both since 6 is the y-iintercept

OpenStudy (anonymous):

I see thank you you've been a great help